Given the following maro-economic model

Y=C+I+G

C=20+0.07Y

I=12+0.1Y

G=10

(a) Express the model in matrix form

(b) Using Cramer's Rule calculate the equilibrium values of the national income (Y), Consumption (C) and Investment (I)

(c) What is the income multiplier?


1
Expert's answer
2021-04-27T07:35:32-0400
"Solution"

"given :\\
Y=C+I+G\\
C=20+0.07Y\\
I=12+0.1Y\\
G=10"


[A] The model in matrix form is:

"Y-C-I=10\\
-0.07Y
+C=20\\
-0.1Y+I=12\\"


"\begin{pmatrix}
 1&-1 & -1\\
 -0.07&1 & 0\\
-0.1&0&1
\end{pmatrix}\begin{pmatrix}
 Y \\C\\I
 
\end{pmatrix}=\begin{pmatrix}
 10\\
 20\\12
\end{pmatrix}"


[B]Using cramer's rule to calculate equilibrium values of Y,C & I:


"\begin{pmatrix}
 1 & -1&-1 \\
 -0.07& 1&0\\
-0.1&0&1
\end{pmatrix}\begin{pmatrix}
 Y \\
 C\\
I
\end{pmatrix}=\begin{pmatrix}
 10 \\
 20\\
12
\end{pmatrix}"


i]Write down the main matrix and find its determinant to be

:"\Delta _0= 0.83"

ii]Replace the 1st column of the main matrix with the solution vector and find its determinant to be: "\Delta_1=42"

iii]Replace the 2nd column of the main matrix with the solution vector and find its determinant to be: "\Delta_2\ =19.54"

iv]Replace the 3rd column of the main matrix with the solution vector and find its determinant to be: "\Delta_3=14.16"

therefore:

"Y=\frac{ Δ1 }{ Δ_0} = \frac{42} 

{0.83} = 50.60\\C= \frac{\Delta_2} 
{ Δ_0} =\frac{ 19.54 }{ 0.83} = 23.54\\I= \frac{Δ_3} { Δ _0}= \frac{14.16 }{ 0.83} = 17.06\\"


[C] income multiplier :

"income \ multiplier[{IM}]=\frac{\Delta GDP_{eq}}{\Delta\ C_{eq}}"


therefore :"\ IM=\frac{50.60}{23.54}=2.15"